OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? A y = 2x + 1 B y = 2x + 3 C y = 2x - 8 D y = 2x - 3

OpenStudy (anonymous):

use this formula y-y1 =m(x-x1) m is slop

OpenStudy (phi):

or just replace x with 2 into a choice, and calculate y. if you get 5, that equation is your answer

OpenStudy (anonymous):

@phi so it would be B?

OpenStudy (phi):

B y = 2x + 3 replace x with 2 y = 2*2 +3 what do you get for y?

OpenStudy (phi):

what is 2 times 2, then add 3 ?

OpenStudy (anonymous):

sorry so its actually A right? @phi

OpenStudy (phi):

A y = 2x+1 replace x with 2 y= 2*2+1 I assume you get 4+1 or 5 so (2,5) is on this line.

OpenStudy (phi):

notice that all the equations have 2x meaning they all have a slope = 2 doing the problem this way does not work if you had 3x (or any number not 2), because then it would have the wrong slope.

OpenStudy (anonymous):

@phi so A is not the answer?

OpenStudy (phi):

If you use maths111 idea, you figure out what the equation is, then look for it. That is the safest way.

OpenStudy (phi):

A is correct. why do you think it isn't ?

OpenStudy (anonymous):

@phi sorry i meant why isn't A the answer, but I'm glad it is can you help me with another?